OpenStudy (anonymous):

1. Set up the following word problem as an equation and solve. Then write your answer as a complete sentence. The ratio of two numbers is 3 to 5. Their sum is 136. Find the two numbers.

OpenStudy (faiqraees):

\[\large\rm \frac{x}{y}= \frac{3}{5}\]\[\large\rm x+y =136 \] Solve it now